OpenStudy (anonymous):

how could you write 8^27 as a product of powers

OpenStudy (anonymous):

Well, 8^27 is multiplying 8 to itself 27 times.

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (anonymous):

@monica12.10 So can you multiply 8 to itself 27 times?

OpenStudy (anonymous):

By "product of powers" I would think the question means to say given something of the form \(a^b=c\), write it as \(c^d=(a^b)^d=a^{bd}\).